OpenStudy (anonymous):

Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of lemon juice are present in the bottle? Show your work. (5 points)

pooja195 (pooja195):

Hey there! Welcome to openstudy! consider reading the CoC: http://openstudy.com/code-of-conduct Heres a link that will help you get to know the site better: http://openstudy.com/study#/updates/548c04e1e4b0f2fbdc23dd6c

OpenStudy (anonymous):

Okay

pooja195 (pooja195):

You're looking for the amout of juice in the bottle and you got: 40% water = 1 liter 60% juice = ? liters 60*1/40= 1.5 liters of juice

OpenStudy (anonymous):

Okay :) Then what

pooja195 (pooja195):

well what do you think?

pooja195 (pooja195):

You could use x : representing number of liters of juice So you'll get 60=40x

OpenStudy (anonymous):

I Came Up with hold on

OpenStudy (anonymous):

0.4 (t) = 1l t -1l over 0.4 = 2.5 ( = 0.6 x 2.5 l

OpenStudy (anonymous):

\[\frac{ 60 }{ 40 }=\frac{ x }{ 1 },x=\frac{ 60 }{ 40}=\frac{ 3 }{ 2 }=1.5~ litre~of~juice.\]

OpenStudy (anonymous):

total 1.5+1=2.5 litres.

OpenStudy (anonymous):

@surjithayer Thats It?

pooja195 (pooja195):

or that could work to :/

OpenStudy (anonymous):

@pooja195 I Have one more i think you can help me figure out :)

pooja195 (pooja195):

sure just report it :)

OpenStudy (anonymous):

In a certain town the temperature, x in degrees Celsius on a certain day is described by two statements: If 3 times the temperature is increased by 2, the temperature is still less than 14°C. Twice the temperature minus 7 is greater than -11°C. Part A: Create a compound inequality to represent the temperature range. (3 points) Part B: Can the temperature in this town be 5°C? Justify your answer by solving the inequalities in Part A. (3 points) Part C: The average temperature in another town is 3°C, but the actual temperature is within 4°C of the average. Write and solve an inequality to find the range of temperature in this town. (4 points)

pooja195 (pooja195):

i meant repost it

OpenStudy (anonymous):

okay hold on